OpenStudy (anonymous):

Working together, it takes two computers 12 minutes to send out a company’s email. If it takes the slower computer 30 minutes to do the job on its own, how long will it take the faster computer to do the job on its own?

OpenStudy (anonymous):

Let x = time (minutes) it takes for the faster computer to do the job alone then 12(1/30 + 1/x) = 1

OpenStudy (anonymous):

multiplying both sides by 30x: 12(x + 30) = 30x 12x + 360 = 30x 360 = 18x 20 minutes = x

OpenStudy (anonymous):

Did that help? :3